OpenStudy (dls):

solve for x: 2 arctan(sinx)=arctan(2secx) x neq to pi/2

OpenStudy (dls):

\[\Huge 2\tan^{−1}(sinx)=\tan^{-1}(2secx)\]

terenzreignz (terenzreignz):

hmm... The one in your post is different from the one ^ up there. Are these two different questions?

OpenStudy (shubhamsrg):

simplify and equate! :O :O

OpenStudy (dls):

same now!

OpenStudy (dls):

\[\LARGE \frac{2sinx}{1+\sin^2x}=\tan^{-1}(\frac{2}{cosx})\]

OpenStudy (shubhamsrg):

re-check LHS :3

OpenStudy (dls):

/??

OpenStudy (dls):

??

OpenStudy (dls):

@shubhamsrg

OpenStudy (shubhamsrg):

denominator will be 1-sin^2x in LHS

OpenStudy (dls):

okay,next?

OpenStudy (shubhamsrg):

its simply after that 1-sin^2x becomes cos^2x then it'll get simplified

OpenStudy (dls):

i am getting tan^2x=1 x=pi/4,-pi/4..so on?

OpenStudy (dls):

can anyone check

OpenStudy (shubhamsrg):

seems right

OpenStudy (dls):

but ans-pi/4

OpenStudy (shubhamsrg):

oh yes,hmm, we'll again put those values back in the original eqn and check which one satisfies

OpenStudy (dls):

no domain given btw

OpenStudy (dls):

I am getting pi/4,-pi/4..3pi/4..-3pi/4..like that.. (-1)^n (2n-1)pi/4..

OpenStudy (dls):

@amistre64 @Callisto @shubhamsrg @experimentX @dan815 @mathslover @terenzreignz

OpenStudy (dan815):

\[\LARGE \frac{2sinx}{cos^{2}x}=\tan^{-1}(\frac{2}{cosx})\] \[\LARGE \frac{2Tan^{2}}{cosx}=(\frac{2}{cosx})\] \[\LARGE \ 2Tan^{2}=({2})\] Tan=1